2. Compare the slope in each of the above equations to the resistance of each resistor.

media%2Ffc8%2Ffc8cc378-3557-44a9-bde0-9a

media%2F5a2%2F5a204272-aa97-411c-bf01-67

i need help with question number 2. I have posted two graphs
Answer #1

From slope, R = 49.8 ohm

given R = 51 ohm

dR = 51-49.8 = 1.2 ohm

dR/R = 1.2 / 51 = 0.0235

% error = 2.35 %